OpenStudy (anonymous):

A board game has a spinner with ten equal sections, numbered 1 through ten. In order to win the game, Mae needs to spin at least a 6 on each of her last two turns. What's the probability of this happening?

OpenStudy (anonymous):

There are 5 numbers that are at least 6 {6,7,8,9,10} There's a 5/10 chance of landing on one of those 5. Spins are independent of one another, so for it to happen twice, that probability would be: (5/10)(5/10) = (1/2)(1/2) = 1/4 or 0.25

OpenStudy (anonymous):

Probabilities of independent events are ALWAYS multiplied. Hope that helps! Take care and don't forget to select a Best Response. Thanks!